OpenStudy (anonymous):

(14,5) (13,2) in a y=mx +b

OpenStudy (anonymous):

The way to approach this problem is to first find the slope and then solve for b. So we begin by finding the slope m= (5-2)/(14-13) which, after a little arithmetic, is m= 3/1 =3. So now we have y=3x+b. Next we need to plug in one of the points into our equation in order to solve for b. So I will choose (14,5). Thus 5=3(14)+b, which equals 5=42+b which equals -37=b. So we have arrived at the solution which is y=3x-37